Sol Summary – April 12th

Crew 178 Daily Summary Report 12 APR 2017
MDRS Daily Summary Report for sol 2
Summary Title: Second EVA, preparations for other experiments and Internet fixing
Mission Status: Good
Sol Activity Summary:
“Today, we again woke up at 8 am after having enjoyed “The Martian” movie yesterday night with the whole team. In the morning, Nathalie invented the “Pancake dance” while she was cooking delicious pancakes for everyone, although her first 5 pancakes were closer to sharcoal than cake. Elke finished the setup of the last parameters for her second EVA using her ground penetrating radar. Calogero continued to configure his software applications for his geomatics EVA. At noon, we ate strange canned turkey meat and a Minestrone bear creek soup with our favorite bread, oven-cooked with love by our Crew Scientist, Mathieu. On the whole, this meal was still a pleasure. In the afternoon, the second EVA was launched under a very sunny weather reaching about 26 degrees Celsius. A fun fact about our Biohazard space suits is that they’re cloaked with an external plastic layer, allowing us to sweat like never inside. Extreme conditions are fulfilled. Quentin and Elke will try to reconstruct a Shepherd’s pie with lyophilized food for an amazing dinner tonight. After dinner, every crew member will have still hard work to finish preparing at best the next days.
Aurian d’Avernas, the Crew Commander”
Look Ahead Plan: The HughesNet technician came today to fix the parable with concrete. We identified that some laptops were automatically burning our daily Internet data with the automatic Windows updates which are very tricky to disable. We finally found the solution and tomorrow will be the first day without any Internet issue. The preparations of the crew experiments are going well. Moreover, we will send our biographies and crew photos tomorrow morning without exception.
Anomalies in work: The wheel of the ground penetrating radar does not work properly. It should be fixed by a phone call to the company of this material tomorrow morning by our Crew Geologist.
Weather: Clear sky and very sunny.
Crew Physical Status: Very good.
EVA: 1 km at the North of the Hab.
Reports to be file: EVA #2 report + Journalist report + Engineer report + Daily Summary Report
Support Requested: none

Sol Summary – April 11th

Summary Title: Technical setups for several experiments and first EVA

Mission Status:  Good

Sol Activity Summary:

We woke up today at 8 am. The Crew Astronomer and the Crew Commander broke the sim for two hours (10-12 am) in order to take the aerial delivery at Hollow Mountain (5 days of late) containing our scientific equipment. We shared meal at 1 pm (tomato soup with bread). The Crew Geologist and the Crew Scientist started their EVA at 2 pm and came back at 3.30 pm. In the meanwhile, the other crew members set up their experiment in the Hab and the Science Lab. The Crew Geologist prepared its plants launched the GreenHab activities. We had to do an emergency EVA (in spacesuits of course) to the Shannon’s trailer from 7 to 9 pm in order to start CapCom session. Indeed, some electronic devices probably consumed all the daily Internet data for this day, that is why we were constrained to exceptionally go to the Shannon’s trailer. At 6.30 pm, we started a group meeting to debrief the day and prepare the day of tomorrow, and also to disable all the electronic devices of each crew member to avoid any further problem with the daily Internet datas. After the evening meal, some of the crew members enjoyed a movie, while others continued to work.Look Ahead Plan: (How are you going to follow up?  What’s the plan for tomorrow?)

Anomalies in work: Internet data were consumed par mégarde

Weather: The beginning of the day started to be dégagé and clear. In the afternoon, some clouds added and the wind became more violent ???

Crew Physical Status: Everyone is fine

EVA: (where did you go?)

Reports to be file: Crew Engineer Report, Journalist Report,

Support Requested: The tunnel structure between the Hab and the Science Lab has suffered from heavy gusts of wind, it will be difficult

Sol Summary – April 10th

Crew 178 Daily Summary Report  10/04/2017

MDRS Daily Summary Report for sol 0

Summary Title: Fixing Internet and aerial delivery delays

Mission Status:  Several technical problems prevent us to start the simulation

Sol Activity Summary:
The Crew Engineer and the Crew Commander started in the morning to fix the Hab’s Internet Connection: it has been a success since we identified that the issue was due to the rotation in the vertical axis of the parable. This rotation varied of a few degrees because of lots of gusts of wind. We then fixed this issue by adding a fastening system with ropes. The rest of the day was dedicated to discover the surrounding areas with ATV’s and rovers to have a better idea of the relief. We also started physical tests (known as the “bip test”) and other physical exercises outside in the framework of the experiment of our physiotherapist. Finally, we passed few hours to fix the aerial delivery delays by calling several customer services. We got the confirmation that our packages will be finally delivered tomorrow at 10 am.

Look Ahead Plan: Internet connection has been fixed and the delivery of our packages should coming tomorrow, we then plan to start the simulation tonight at 6 pm although two of our crew members will have to break the sim tomorrow morning to get our aerial delivery.

Anomalies in work:  The technical assistance of HughesNet should still come to install a better fastening system for the parable since our solution is not robust against strong winds.

Weather: The weather was cleared and the sun predominant.

Crew Physical Status: Perfect health


Reports to be file: Crew Engineer Report + Daily Summary

Support Requested:

Sol Summary – March 31st

MDRS Daily Summary Report for SOL 6, 31Mar2017
Person filing out Report: Victoria LaBarre
Position: Crew OX and Crew Chemist
Summary Title: Clean up
Mission Status: Concluding experiments and EVAs and cleaning Hab for next crew
Sol Activity Summary: Packed up robot, clock and bike generator, Microbiologist EVA, cleaned entire Hab
Look-Ahead Plan: request ability or help in handling toilet issue/End simulation and go back to Grand Junction.
Anomalies in work:  Nothing out of the ordinary
-Morning: Overcast, windy
-Afternoon-Sunny, windy, showers
Crew Physical Status: Crew is healthy

Sol Summary – March 29th

Crew 177 Sol 4 Summary 29Mar2017
Person filling out Report:
Pitchayapa Jingjit—a biologist and a photographer for Crew 177
Summary Title:
Sol 4 at Mars
Mission Status:
Everyone is doing great and is hopeful about the toilet situation
SOL Activity Summary:
Improving individual/group research project and continuing the chemistry EVA. Depending on tonight’s sky, astronomy EVA may be conducted.
Look ahead Plan:
Continuing the geology EVA at the Little Canyon and depending on tonight’s sky, the astronomy EVA may be performed, and progressing through with individual/group projects.
Anomalies in Progress:
The weather is sunny in the morning and afternoon. We are expecting a clear sky tonight.
Crew Physical Status:
Everyone is better adjusted at Mars and is happy that the plumber came to the HAB.


Sol Summary – March 27th

Person filling out Report:
Elijah Espinoza

Summary Title:
First full day on Mars

Mission Status:
Troubleshooting Research Projects

SOL Activity Summary:
Rover related tasks, charging bike tasks, EVA, Settling in

Look ahead Plan:
EVA, Reporter coming tomorrow

Anomalies in Progress:

Cloudy and a little rain in the morning, sunny and partly cloudy after noon

Crew Physical Status:
Everyone is healthy and feels great! Morale is high!

Reports to be Filled:

-Operations Report

-Commanders Report

-EVA Report

-EVA Request

Sol Summary – March 26th

MDRS Daily Summary Report for SOL 1

Person filing out Report: Esteban Ramirez

Summary Title: Crew 177 Enters Mars

Mission Status: Hard at work

Sol Activity Summary: Unpacking, organizing supplies, EVA, generator related tasks, rover related tasks

Look Ahead Plan: EVA, handle water issue

Anomalies in work:  Nothing out of the ordinary

Weather: Sunny, clear skies

Crew Physical Status: Crew is healthy

Reports to be filled:

-Operations Report

-Commanders Report

-EVA Report

-EVA Request

Sol Summary – March 25th

Commander Report Crew 177:
Otsmar Villarroel
Sol 1
Earth date: 03/25/2017

Good evening!

Today we met the very helpful Crew 176 at MDRS.  Upon arrival at the hab, crew members were trained by Crew 176 members.  Training included a brief tour around the MDRS facilities as well as over use of the ATVs, operation report data, water leak issue, etc.  After the training we went to Bull Mountain Market to pick up our food according to the inventory sent by crew 176 as requested by Shannon.

We’ve made a plan for tomorrow, which will include, pending approval, our Orientation EVAs. We are expecting to enter sim tomorrow morning at 9:30 am, if possible.

From Crew 177; Hanksville, UT
Crew Commander

Sol Summary – March 24th

MDRS Daily Summary Report for SOL 13

Person filling out Report: Michał Kazaniecki

Summary Title: A step closer to the successful end

Mission Status: Busy, busy, busy

Sol Activity Summary: EVA,  rover related tasks, cleaning

Look Ahead Plan: more cleaning and packing

Anomalies in work: everything as expected

Weather: sunny, but not too hot

Crew Physical Status: everyone feels good

Reports to be filled:

-Operations Report
-Commander Report
-EVA Report
-EVA Request
-Summary Report (as requested by Robert Zubrin)

Sol Summary – March 22nd

MDRS Daily Summary Report for SOL 11 (22Mar2017)

Person filling out Report: Michał Kazaniecki

Summary Title: Internet issues and beyond 

Mission Status: Still work in progress

Sol Activity Summary: EVA, Internet “fixing”, rover testing, installing the experimental shower

Look Ahead Plan: EVA with rover, tests of rover Ares

Anomalies in work: Unstable Internet connection (again)

Weather: Cloudy and windy, heavy rain and strong wind at night

Crew Physical Status: one minor toe injury, everyone else feels good

Reports to be filled:

-Operations Report
-Commander Report
-EVA Report
-EVA Request (via SMS)
-Science Report